# Pebble Games and Algebraic Proof Systems

### v1

@@
Abstract
Analyzing refutations of the well known pebbling formulas $\mathrm{Peb}(G)$ we prove some new strong connections between pebble games and algebraic proof system, showing that there is a parallelism between the reversible, black and black-white pebbling games on one side, and the three algebraic proof systems Nullstellensatz, Monomial Calculus and Polynomial Calculus on the other side. In particular we prove that for any DAG $G$ with a single sink, if there is a Monomial Calculus refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $t$ then there is a black pebbling strategy on $G$ with space $s$ and time $t+s$. Also if there is a black pebbling strategy for $G$ with space $s$ and time $t$ it is possible to extract from it a MC refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $ts$. These results are analogous to those proven in [dRMNR21] for the case of reversible pebbling and Nullstellensatz. Using them we prove degree separations between NS, MC and PC, as well as strong degree-size tradeoffs for $\mathrm{MC}$.
We also notice that for any directed acyclic graph $G$ the space needed in a pebbling strategy on $G$, for the three versions of the game, reversible, black and black-white, exactly matches the variable space complexity of a refutation of the corresponding pebbling formula $\mathrm{Peb}(G)$ in each of the algebraic proof systems $\mathrm{NS},\mathrm{MC}$ and $\mathrm{PC}$. Using known pebbling bounds on graphs, this connection implies separations between the corresponding variable space measures.
1 Introduction
The use of pebble games in complexity theory goes back many decades. They offer a very clean tool to analyze certain complexity measures, mainly space and time, in an isolated way on a graph, which can then be translated to specific computational models. Very good overviews of these results can be found in [Pip80, Sav98, Nor15].
@@ 1 Introduction
In proof complexity one tries to understand the resources needed for a proof of a mathematical statement in a formalized system. Pebbling games have also become one of the most useful tools for proving results in this area. The reason for this is that one can often translate a certain measure for the pebbling game, mainly number of pebbles or pebbling time, into a suitable complexity measure for a concrete proof system. Very often the bounds for this measure in a graph translate accurately to complexity bounds in the different proof systems for a certain kind of contradictory formulas mimicking the game, called pebbling formulas. These formulas were introduced in [BW01] and have been extremely useful for proving separations, upper and lower bounds as well as tradeoff results in basically all studied proof systems. See e.g. [Nor13].
@@
In the present paper we will concentrate on algebraic proof system. In these systems formulas are encoded as sets of polynomials over a field and the question of whether a formula is unsatisfiable is translated to the question of whether the polynomials have a common root. Powerful algebraic tools like the Gröbner Basis Algorithm can be used for this purpose. Several algebraic proof systems have been introduced in the literature (defined formally bellow). Well known are Nullstellensatz (NS) introduced in [BIK+94] and the more powerful Polynomial Calculus (PC) defined in [CEI96]. The first one is usually considered as a static system in which a “one-shot” proof has to be produced, while in PC there are certain derivation rules like in a more standard proof system.
@@
The best studied complexity measures for refutations in these systems are the degree (maximum degree of a polynomial in the refutation) and size (number of monomials in the refutation, counted with repetitions). For studying the connections with the pebble games it is very useful to consider also space measures and the configurational refutations associated with space. We will use the variable space measure (number of variables that are simultaneously active in a refutation).
@@
In [BG15] the Monomial Calculus system (MC) was identified. This system is defined by limiting the multiplication rule in PC to monomials and its power lies between $\mathrm{NS}$ and $\mathrm{PC}$. Building on results from [AM13] for the Sherali-Adams proof system, the authors proved that for any pair of non-isomorphic graphs, the MC degree for the refutation of the corresponding isomorphism formulas exactly corresponds to the Weisfeiler-Leman bound for separating the graphs, a very important tool in graph theory and descriptive complexity. This equivalence (as well as the relations to pebbling shown here) motivates the study of Monomial Calculus as a natural proof system between NC and PC.
@@
As mentioned above, connections between pebbling games and algebraic systems have been known since [BCIP02]. It was proved in the mentioned reference that for any directed acyclic graph (DAG) $G$ the corresponding pebbling formula $\mathrm{Peb}(G)$ can be refuted with constant degree in PC but it requires degree $\Omega(s)$ in NS, where $s$ is the black pebbling price of $G,\mathsf{Black}(G)$. Using pebbling results, this automatically proves a strong degree separation between NS and PC. As a more recent example, the authors in [dRMNR21] proved a very tight connection between NS and the reversible pebbling game. They showed that space and time in the game played on a DAG exactly correspond to the degree and size measures in a NS refutation of the corresponding pebbling formula. From this connection strong degree-size tradeoffs for NS follow. This result also improves the degree separation from [BCIP02] since it is known that there are graphs for which the reversible pebbling price is a logarithmic factor larger than the black pebbling price.
@@ 1.1 Our Results
In Section 3 we prove that very similar results to those given in [dRMNR21] for NS and reversible pebbling are also true for the case of MC and black pebbling. More concretely we show in Theorem 14 that for any DAG $G$ with a single sink, if there is a MC refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $t$ then there is a black pebbling strategy on $G$ with space $s$ and time $t+s$. This is done by proving that any Horn formula has a very especial kind of MC refutation, which we call input MC refutation since it is the same concept as an input refutation in Resolution. Horn formulas constitute an important class with applications in many areas like program verification or logic programming. It is well known that input Resolution is complete for Horn formulas.
@@
For the other direction, we show in Theorem 9 that from a black pebbling strategy for $G$ with space $s$ and time $t$ it is possible to extract a MC refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $ts$. The small loss in the time parameter compared to the results in [dRMNR21] comes from the fact that size complexity is measured in slightly different ways in NS and MC. Using these results we are able to show degree separations between NS and MC as well as the first strong degree separations between MC and PC. We also use the simultaneous relation with time and space in the black pebbling game to obtain strong degree-size tradeoffs for MC in the same spirit as those in [dRMNR21]. These results show that there are families of unsatisfiable formulas that have polynomial size MC refutations with a certain degree, but the size of the refutation becomes superpolynomial if the degree decreases. The results also show that strong degree lower bounds for MC refutations do not imply exponential size lower bounds as it happens in the PC proof system [IPS99].
@@
From the results in [dRMNR21] and this paper follows that the degrees of the refutations of pebbling formulas in NS and MS correspond exactly to the space in reversible and black pebbling games respectively. It would be a nice result if the same could be said about PC degree and space in the black-white game. Unfortunately this is not the case since as mentioned above, it was proven in [BCIP02] that for any DAG the corresponding pebbling formula can be refuted within constant PC degree. We notice however that if instead of the degree we consider the complexity measure of variable space, then the connection still holds. We notice that for for any single sink DAG $G$ the variable space complexity of refuting $\mathrm{Peb}(G)$ in each of the algebraic proof systems $\mathrm{NS},\mathrm{MC}$ and $\mathrm{PC}$ is exactly the variable space needed in a strategy for pebbling $G$ in each of the three versions Reversible, Black and Black-White of the pebble game. These results allow us to apply known separations between the pebbling space needed in the different versions of the the game, in order to obtain the first separations in the variable space measure between the different proof systems.
2 Preliminaries
2.1 Pebble Games
@@
Black pebbling was first mentioned implicitly in [PH70], while black-white pebbling was introduced in [CS76]. Note, that there exist several variants of the (black-white) pebble game in the literature. For differences between these variants, we refer to [Nor15]. For the following definitions, let $G=(V,E)$ be a DAG with a unique sink vertex $z$.
Definition 1 (Black and black-white pebble games).
@@
The black-white pebble game on $G$ is the following one-player game: At any time $i$ of the game, there is a pebble configuration $\mathbb{P}_{i}:=(B_{i},W_{i})$, where $B_{i}\cap W_{i}=\varnothing$ and $B_{i}\subseteq V$ is the set of black pebbles and $W_{i}\subseteq V$ is the set of white pebbles, respectively. A pebble configuration $\mathbb{P}_{i-1}=(B_{i-1},W_{i-1})$ can be changed to $\mathbb{P}_{i}=(B_{i},W_{i})$ by applying exactly one of the following rules:
Black pebble placement on $v$:
@@
Black pebble removal from $v$:
@@
White pebble placement on $v$:
@@
White pebble removal from $v$:
@@
Definition 2 (Pebbling time, space, and price).
The time of a pebbling $\mathcal{P}=(\mathbb{P}_{0},\mathbb{P}_{1},\dots,\mathbb{P}_{t})$ is $\mathsf{time}(\mathcal{P}):=t$ and the space of it is $\mathsf{space}(\mathcal{P}):=\max_{i\in[t]}|B_{i}\cup W_{i}|$. The black-white pebbling price (also known as the pebbling measure or pebbling number) of $G$, which we will denote by $\mathsf{BW}(G)$, is the minimum space of any black-white pebbling of $G$. The black pebbling price of $G$, denoted by $\mathsf{Black}(G)$, is the minimum space of any black pebbling of $G$. By the observation above, the white pebbling price $\mathsf{White}(G)$ coincides with $\mathsf{Black}(G)$
Finally, we mention the reversible pebble game introduced in [Ben89]. In the reversible pebble game only black pebbles are placed and the moves performed in reverse order should also constitute a legal black pebbling, which means that the rules for pebble placements and removals have to become symmetric.
Definition 3 (Reversible pebble game).
The reversible pebble game on $G$ is the following version of the game: At any time $i$ of the game, we have a pebble configuration $\mathbb{P}_{i}\subseteq V$. A configuration $\mathbb{P}_{i-1}$ can be changed to $\mathbb{P}_{i}$ by applying exactly one of the following rules:
Pebble placement on $v$:
@@
Reversible pebble removal from $v$:
@@ 2.2 Formulas and Polynomials
A Horn formula in a special type of CNF formula in which each clause has at most one positive literal. For a more detailed treatment of formulas as well as the well known Resolution proof system we refer the interested reader to some of the introductory texts in the area like [ST13]. We will basically only deal with pebbling formulas. These provide the connection between pebbling games and proof complexity.
Definition 4 (Pebbling formulas).
Let $G=(V,E)$ be a DAG with a set of sources $S\subseteq V$ and a unique sink $z$. We identify every vertex $v\in V$ with a Boolean variable $x_{v}$. For a vertex $v\in V$ we denote by $\operatorname{\mathrm{p}red}(v)$ the set of its direct predecessors. In particular, for a source vertex $v$, $\operatorname{\mathrm{p}red}(v)=\emptyset$. The pebbling contradiction over $G$, denoted $\mathrm{Peb}(G)$, is the conjunction of the following clauses:
@@
A way to prove that a CNF formula is unsatisfiable is by translating it into a set of polynomials over a field $\mathbb{F}$ and then show that these polynomials do not have any common $\{0,1\}$-valued root. A clause $C=\bigvee_{x\in P}x\vee\bigvee_{y\in N}\bar{y}$ can be encoded as the polynomial $p(C)=\prod_{x\in P}(1-x)\prod_{y\in N}y$. A set of clauses $C_{1},\dots,C_{m}$ is translated as set of polynomials $p(C_{1}),\dots,p(C_{m})$. Adding the polynomials $x_{i}^{2}-x_{i}$ (as axioms) for each variable $x_{i}$, there is no common $\{0,1\}$-valued root for all these polynomials if and only if the original set of clauses is unsatisfiable. The intuition here is to identify false with 1 and true with 0. A monomial is falsified by a Boolean assignment if all its variables get value 1, while it is satisfied if one of its variables gets value 0. In this context we will consider a monomial $m$ as a set of variables and a polynomial $p$ as a linear combination of monomials. A monomial with its coefficient in $\mathbb{F}$ is called a monomial term.
@@
2.3 Algebraic Proof Systems
Several proof systems that work with polynomials have been defined in the literature. The simplest one is Nullstellensatz, $\mathrm{NS}$.
Definition 5.
@@ 2.3 Algebraic Proof Systems
A refutation in $\mathrm{PC}$ of an initial unsolvable set of polynomials ${\cal P}$ is a sequence of polynomials $\{q_{1},\dots,q_{m}\}$ such that each $q_{i}$ is either a polynomial in ${\cal P}$, a Boolean axiom ${x^{2}_{i}-x_{i}}$ or it is obtained by previous polynomials in the sequence applying one of the rules of the calculus.
A third less known algebraic proof system between $\mathrm{NS}$ and $\mathrm{PC}$ is Monomial Calculus, $\mathrm{MC}.$ This system was introduced in [BG15] identifying exactly the complexity of refuting graph isomorphism formulas. This proof system is defined like $\mathrm{PC}$ but the multiplication rule is only allowed to be applied to a monomial, or to a monomial times an axiom.
Definition 7.
@@
As is the case of $\mathrm{PC},$ a refutation in $\mathrm{MC}$ of an initial unsolvable set of polynomials ${\cal P}$ is a sequence of polynomials $\{q_{1},\dots,q_{m}\}$ where each one of them is either in ${\cal P}$, an axiom or is obtained by applying one of the rules of the calculus.
As pointed out in [BG15], an equivalent definition of the Nullstellensatz system, but a dynamic one, would be to restrict the multiplication rule in the above definition even more, and only allow to apply it to polynomials that are a monomial multiplied by an axiom. In this way, the difference in the definition of the three systems NS, MC and PC is just a variation on how the multiplication rule can be applied: multiplying a variable $x_{i}$ by any polynomial in the PC case, by a monomial or by a monomial times an axiom in the MC case and only by a monomial times an axiom in the NS case. This alternative view of the definition also allows to consider configurational proofs in the $\mathrm{NS}$ system. In order to analyze and compare refutations we will consider several complexity measures on them.
Definition 8 (Complexity measures).
Let ${\cal C}$ be one of the mentioned systems ${\cal C}\in\{\mathrm{NS},\mathrm{MC},\mathrm{PC}\}$ and let ${\pi}=\{q_{1},\dots,q_{m}\}$ be a ${\cal C}$ refutation. The degree of a polynomial $q_{i}$, $\deg(q_{i}),$ is the maximum degree of its monomials and the degree of ${\pi}$, $\deg_{\cal C}(\pi)=\max_{i=1,\ldots,n}(\deg(q_{i}))$. The size of ${\pi}$, denoted by $\mathrm{Size}_{\cal C}({\pi})$ is the total number of monomials in $\pi$ (counted with repetitions), when all polynomials $p_{i}$ are fully expanded as linear combinations of monomials [^1] Usually the size in the $\mathrm{NS}$ proof system is defined in a different way, for simplicity we keep this unifying definition although in some of the referenced results the size of $\mathrm{NS}$ refutations corresponds to the size definition given in the reference..
For the space measures we need to define configurational proofs. Such a proof $\pi$ of a set of polynomials ${\cal P}$ in the system ${\cal C}$ is a sequence of configurations $\pi=C_{0},\dots C_{t}$ in which each $C_{i}$ is a set of polynomials with $C_{0}=\emptyset$ and $C_{t}=\{1\}$. Each configuration $C_{i}$ represents a set of polynomials that are kept simultaneously in memory at time $i$ in the refutation, and for each $i,0<i\leq t,$ $C_{i}$ is either
@@
  • $C_{i-1}\cup\{p\}$ for some $p$ inferred from $C_{i-1}$ and ${\cal P}$ by some rule of the system ${\cal C}$ (inference).
The variable space of the proof $\pi$, $\mathrm{VSpace}_{\cal C}({\pi})$ is defined as the maximum number of different variables appearing in any configuration of the proof.
@@
3 Monomial Calculus and pebbling formulas
In [dRMNR21] it was shown that for any DAG $G$ with a single sink, the reversible pebbling space and time of $G$, exactly coincides with the degree and the size of a NS refutation of $\mathrm{Peb}_{G}$. We show that a very similar relation holds for the case of black pebbling and Monomial Calculus.

### v2

@@
Abstract.
Analyzing refutations of the well known pebbling formulas $\mathrm{Peb}(G)$ we prove some new strong connections between pebble games and algebraic proof system, showing that there is a parallelism between the reversible, black and black-white pebbling games on one side, and the three algebraic proof systems Nullstellensatz, Monomial Calculus and Polynomial Calculus on the other side. In particular we prove that for any directed acyclic graph $G$ with a single sink, if there is a Monomial Calculus refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $t$ then there is a black pebbling strategy on $G$ with space $s$ and time $t+s$. Also if there is a black pebbling strategy for $G$ with space $s$ and time $t$ it is possible to extract from it a MC refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $ts$. These results are analogous to those proven in [15] for the case of reversible pebbling and Nullstellensatz. Using them we prove degree separations between Nullstellensatz, Monomial Calculus and Polynomial Calculus, as well as strong degree-size tradeoffs for Monomial Calculus.
We also notice that for any directed acyclic graph $G$ the space needed in a pebbling strategy on $G$, for the three versions of the game, reversible, black and black-white, exactly matches the variable space complexity of a refutation of the corresponding pebbling formula $\mathrm{Peb}(G)$ in each of the algebraic proof systems Nullstellensatz, Monomial Calculus and Polynomial Calculus. Using known pebbling bounds on graphs, this connection implies separations between the corresponding variable space measures.
Key words and phrases:
proof complexity, pebble games, algebraic proof systems
1. Introduction
The use of pebble games in complexity theory goes back many decades. They offer a very clean tool to analyze certain complexity measures, mainly space and time, in an isolated way on a graph, which can then be translated to specific computational models. Very good overviews of these results can be found in [29, 31, 26].
@@ 1. Introduction
In proof complexity one tries to understand the resources needed for a proof of a mathematical statement in a formalized system. Pebbling games have also become one of the most useful tools for proving results in this area, as can be seen for example in [23],[5], [24], [25], [33], [15], to name just a few references. The reason for this is that one can often translate a certain measure for the pebbling game, mainly number of pebbles or pebbling time, into a suitable complexity measure for a concrete proof system. Very often the bounds for this measure in a graph translate accurately to complexity bounds in the different proof systems for a certain kind of contradictory formulas mimicking the game, called pebbling formulas. These formulas were introduced in [6] and have been extremely useful for proving separations, upper and lower bounds as well as tradeoff results in basically all studied proof systems. See e.g. [25].
@@
In the present paper we will concentrate on algebraic proof system. In these systems formulas are encoded as sets of polynomials over a field and the question of whether a formula is unsatisfiable is translated to the question of whether the polynomials have a common root. Powerful algebraic tools like the Gröbner Basis Algorithm can be used for this purpose. Several algebraic proof systems have been introduced in the literature (defined formally bellow). Well known ones are Nullstellensatz (NS) introduced in [3] and the more powerful Polynomial Calculus (PC) defined in [12]. The first one is usually considered as a static system in which a “one-shot” proof has to be produced, while in PC there are certain derivation rules like in a more standard proof system.
@@
The best studied complexity measures for refutations in these systems are the degree (maximum degree of a polynomial in the refutation) and size (number of monomials in the refutation, counted with repetitions). For studying the connections with the pebble games it is very useful to consider also space measures and the configurational refutations associated with space. We will use the variable space measure (number of variables that appear simultaneously in a refutation).
@@
In [9] the Monomial Calculus system (MC) was identified. This system is defined by limiting the multiplication rule in PC to monomials, and its power lies between $\mathrm{NS}$ and $\mathrm{PC}$. Building on results from [2] for the Sherali-Adams proof system, the authors proved that for any pair of non-isomorphic graphs, the MC degree for the refutation of the corresponding isomorphism formulas exactly corresponds to the Weisfeiler-Leman bound for separating the graphs, a very important tool in graph theory and descriptive complexity. This equivalence (as well as the relations to pebbling shown here) motivates the study of Monomial Calculus as a natural proof system between NC and PC.
@@
As mentioned above, connections between pebbling games and algebraic systems have been known since [10]. It was proved in the mentioned reference that for any directed acyclic graph (DAG) $G$ the corresponding pebbling formula $\mathrm{Peb}(G)$ can be refuted with constant degree in PC but it requires degree $\Omega(s)$ in NS, where $s$ is the black pebbling price of $G,\mathsf{Black}(G)$. Using pebbling results, this automatically proves a strong degree separation between NS and PC. As a more recent example, the authors in [15] proved a very tight connection between NS and the reversible pebbling game. They showed that space and time in the game played on a DAG exactly correspond to the degree and size measures in a NS refutation of the corresponding pebbling formula. From this connection strong degree-size tradeoffs for NS follow. This result also improves the degree separation from [10] since it is known that there are graphs for which the reversible pebbling price is a logarithmic factor larger than the black pebbling price.
@@ 1.1. Our Results
In Section 3 we prove that very similar results to those given in [15] for NS and reversible pebbling are also true for the case of MC and black pebbling. More concretely we show in Theorem 4 that for any DAG $G$ with a single sink, if there is a MC refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $t$ then there is a black pebbling strategy on $G$ with space $s$ and time $t+s$. This is done by proving that any Horn formula has a very especial kind of MC refutation, which we call input MC refutation since it is the same concept as an input refutation in Resolution. Horn formulas constitute an important class with applications in many areas like program verification or logic programming. It is well known that input Resolution is complete for Horn formulas.
@@
For the other direction, we show in Theorem 1 that from a black pebbling strategy for $G$ with space $s$ and time $t$ it is possible to extract a MC refutation for $\mathrm{Peb}(G)$ having simultaneously degree $s$ and size $ts$. The small loss in the time parameter compared to the results in [15] comes from the fact that size complexity is measured in slightly different ways in NS and MC. Using these results we are able to show degree separations between NS and MC as well as the first strong degree separations between MC and PC. We also use the simultaneous relation with time and space in the black pebbling game to obtain strong degree-size tradeoffs for MC in the same spirit as those in [15]. These results show that there are families of unsatisfiable formulas that have polynomial size MC refutations with a certain degree, but the size of the refutation becomes superpolynomial if the degree decreases. The results also show that strong degree lower bounds for MC refutations do not imply exponential size lower bounds as happens in the PC proof system [20].
@@
From the results in [15] and this paper follows that the degrees of the refutations of pebbling formulas in NS and MS correspond exactly to the space in reversible and black pebbling games respectively. It would be a nice result if the same could be said about PC degree and space in the black-white game. Unfortunately this is not the case since as mentioned above, it was proven in [10] that for any DAG the corresponding pebbling formula can be refuted within constant PC degree. We notice however that if instead of the degree we consider the complexity measure of variable space, then the connection still holds. We notice that for any single sink DAG $G$ the variable space complexity of refuting $\mathrm{Peb}(G)$ in each of the algebraic proof systems $\mathrm{NS},\mathrm{MC}$ and $\mathrm{PC}$ is exactly the space needed in a strategy for pebbling $G$ in each of the three versions Reversible, Black and Black-White of the pebble game. These results allow us to apply known separations between the pebbling space needed in the different versions of the game, in order to obtain the first separations in the variable space measure between the different proof systems.
2. Preliminaries
2.1. Pebble Games
@@
Black pebbling was first mentioned implicitly in [27], while black-white pebbling was introduced in [13]. Note that there exist several variants of the (black-white) pebble game in the literature. For differences between these variants, we refer to [26]. For the following definitions, let $G=(V,E)$ be a DAG with a unique sink vertex $z$.
{defi}
@@
[Black and black-white pebble games] The black-white pebble game on $G$ is the following one-player game: At any time $i$ of the game, there is a pebble configuration $\mathbb{P}_{i}:=(B_{i},W_{i})$, where $B_{i}\cap W_{i}=\varnothing$ and $B_{i}\subseteq V$ is the set of black pebbles and $W_{i}\subseteq V$ is the set of white pebbles, respectively. A pebble configuration $\mathbb{P}_{i-1}=(B_{i-1},W_{i-1})$ can be changed to $\mathbb{P}_{i}=(B_{i},W_{i})$ by applying exactly one of the following rules:
Black pebble placement on $v$::
@@
Black pebble removal from $v$::
@@
White pebble placement on $v$::
@@
White pebble removal from $v$::
@@
{defi}
[Pebbling time, space, and price] The time of a pebbling $\mathcal{P}=(\mathbb{P}_{0},\mathbb{P}_{1},\dots,\mathbb{P}_{t})$ is $\mathsf{time}(\mathcal{P}):=t$ and its space is $\mathsf{space}(\mathcal{P}):=\max_{i\in{\rm Nature}{t}}|B_{i}\cup W_{i}|$. The black-white pebbling price (also known as the pebbling measure or pebbling number) of $G$, which we will denote by $\mathsf{BW}(G)$, is the minimum space of any black-white pebbling of $G$. The black pebbling price of $G$, denoted by $\mathsf{Black}(G)$, is the minimum space of any black pebbling of $G$. By the observation above, the white pebbling price $\mathsf{White}(G)$ coincides with $\mathsf{Black}(G)$
Finally, we mention the reversible pebble game introduced in [8]. In the reversible pebble game only black pebbles are placed and the moves performed in reverse order should also constitute a legal black pebbling, which means that the rules for pebble placements and removals have to become symmetric.
{defi}
[Reversible pebble game] The reversible pebble game on $G$ is the following version of the game: At any time $i$ of the game, we have a pebble configuration $\mathbb{P}_{i}\subseteq V$. A configuration $\mathbb{P}_{i-1}$ can be changed to $\mathbb{P}_{i}$ by applying exactly one of the following rules:
Pebble placement on $v$::
@@
Reversible pebble removal from $v$::
@@ 2.2. Formulas and Polynomials
A Horn formula in a special type of CNF formula in which each clause has at most one positive literal. For a more detailed treatment of formulas as well as the well known Resolution proof system we refer the interested reader to some of the introductory texts in the area like [32]. We will basically only deal with pebbling formulas. These provide the connection between pebbling games and proof complexity.
{defi}
[Pebbling formulas] Let $G=(V,E)$ be a DAG with a set of sources $S\subseteq V$ and a unique sink $z$. We identify every vertex $v\in V$ with a Boolean variable $x_{v}$. For a vertex $v\in V$ we denote by $\operatorname{\mathrm{p}red}(v)$ the set of its direct predecessors. In particular, for a source vertex $v$, $\operatorname{\mathrm{p}red}(v)=\emptyset$. The pebbling contradiction over $G$, denoted $\mathrm{Peb}(G)$, is the conjunction of the following clauses:
@@
A way to prove that a CNF formula is unsatisfiable is by translating it into a set of polynomials over a field $\mathbb{F}$ and then show that these polynomials do not have any common $\{0,1\}$-valued root. A clause $C=\bigvee_{x\in P}x\vee\bigvee_{y\in N}\bar{y}$ can be encoded as the polynomial $p(C)=\prod_{x\in P}(1-x)\prod_{y\in N}y$. A set of clauses $C_{1},\dots,C_{m}$ is translated as set of polynomials $p(C_{1}),\dots,p(C_{m})$. Adding the polynomials $x_{i}^{2}-x_{i}$ (as axioms) for each variable $x_{i}$, there is no common $\{0,1\}$-valued root for all these polynomials if and only if the original set of clauses is unsatisfiable. In this context we will consider a monomial $m$ as a set of variables and a polynomial $p$ as a linear combination of monomials. A monomial with its coefficient in $\mathbb{F}$ is called a monomial term.
@@
2.3. Algebraic Proof Systems
Several proof systems that work with polynomials have been defined in the literature. The simplest one is Nullstellensatz, $\mathrm{NS}$. {defi}
@@ 2.3. Algebraic Proof Systems
A refutation in $\mathrm{PC}$ of an initial unsolvable set of polynomials $\mathcal{P}$} is a sequence of polynomials $\{q_{1},\dots,q_{m}\}$ such that each $q_{i}$ is either a polynomial in $\mathcal{P}$}, a Boolean axiom $x^{2}_{i}-x_{i}$} or it is obtained by previous polynomials in the sequence applying one of the rules of the calculus.
A third less known algebraic proof system between $\mathrm{NS}$ and $\mathrm{PC}$ is Monomial Calculus, $\mathrm{MC}.$ This system was introduced in [9] identifying exactly the complexity of refuting graph isomorphism formulas. This proof system is defined like $\mathrm{PC}$ but the multiplication rule is only allowed to be applied to a monomial, or to a monomial times an axiom.
{defi}
@@
As is the case of $\mathrm{PC},$ a refutation in $\mathrm{MC}$ of an initial unsolvable set of polynomials $\mathcal{P}$} is a sequence of polynomials $\{q_{1},\dots,q_{m}\}$ where each one of them is either in $\mathcal{P}$}, an axiom or is obtained by applying one of the rules of the calculus.
As pointed out in [9], an equivalent definition of the Nullstellensatz system, but a dynamic one, would be to restrict the multiplication rule in the above definition even more, and only allow to apply it to polynomials that are a monomial multiplied by an axiom. In this way, the difference in the definition of the three systems NS, MC and PC is just a variation on how the multiplication rule can be applied: multiplying a variable $x_{i}$ by any polynomial in the PC case, by a monomial or by a monomial times an axiom in the MC case and only by a monomial times an axiom in the NS case. This alternative view of the definition also allows to consider configurational proofs in the $\mathrm{NS}$ system.
{defi}
[Configurational proofs] Let $\mathcal{C}$} be one of the mentioned systems ${\mathcal{C}}\in\{\mathrm{NS},\mathrm{MC},\mathrm{PC}\}$. For the space measures we need to define configurational proofs. Such a proof $\pi$ of a set of polynomials $\mathcal{P}$} in the system $\mathcal{C}$} is a sequence of configurations $\pi=C_{0},\dots C_{t}$ in which each $C_{i}$ is a set of polynomials with $C_{0}=\emptyset$ and $C_{t}=\{1\}$. Each configuration $C_{i}$ represents a set of polynomials that are kept simultaneously in memory at time $i$ in the refutation, and for each $i,0<i\leq t,$ $C_{i}$ is either
@@
$C_{i-1}\cup\{p\}$ for some $p$ inferred from $C_{i-1}$ and $\mathcal{P}$} by some rule of the system $\mathcal{C}$} (inference).
In order to analyze and compare refutations we will consider several complexity measures on them.
{defi}
[Complexity measures] Let $\mathcal{C}$} be one of the mentioned systems ${\mathcal{C}}\in\{\mathrm{NS},\mathrm{MC},\mathrm{PC}\}$ and let ${\pi}=\{q_{1},\dots,q_{m}\}$ be a $\mathcal{C}$} refutation. The degree of a polynomial $q_{i}$, $\deg(q_{i}),$ is the maximum degree of its monomials and the degree of $\pi$}, $\deg_{\mathcal{C}}(\pi)=\max_{i=1,\ldots,m}(\deg(q_{i}))$. The size of $\pi$}, denoted by $\mathrm{Size}_{\mathcal{C}}({\pi})$ is the total number of monomials in $\pi$ (counted with repetitions), when all polynomials $p_{i}$ are fully expanded as linear combinations of monomials [^1] Usually the size in the $\mathrm{NS}$ proof system is defined in a different way, for simplicity we keep this unifying definition although in some of the referenced results the size of $\mathrm{NS}$ refutations corresponds to the size definition given in the reference..
The variable space of the proof $\pi$, $\mathrm{VSpace}_{\mathcal{C}}({\pi})$ is defined as the maximum number of different variables appearing in any configuration of the proof.
@@
3. Monomial Calculus and pebbling formulas
In [15] it was shown that for any DAG $G$ with a single sink, the reversible pebbling space and time of $G$, exactly coincides with the degree and the size of a NS refutation of $\mathrm{Peb}_{G}$. We show that a very similar relation holds for the case of black pebbling and Monomial Calculus.